how do you graph the system y= -2x - 1??
Answer by Fombitz(32388)

Find two points.
First let x=0, y=-2(0)-1=-1, graph (0,-1).
Then let y=0, 0=-2x-1, 2x=-1, x=-1/2, graph (-1/2,0).
Connect the points with the line.